# Structure-Based Molecule Optimization via Gradient-Guided Bayesian Update

### v2

@@ 5.2 Unconstrained Optimization
The optimized distribution greatly improves upon the original generated distribution, as shown in the distribution shift in Figure 3 for single objective optimization, and Table 1 (row 13 vs. row 9).
As shown in Table 1, our model achieves state-of-the-art in affinity-related metrics while being highly drug-like, with the best Success Rate of 51.3%, a fourthfold improvement over TAGMol (row 13 vs. row 12).
RGA (Fu et al., 2022) and DecompOpt (Zhou et al., 2024) show satisfactory Success Rate, enjoying the advantage of oracle-based screening at some expense of diversity. Given the same use of Z-score (Zhou et al., 2024), we report a variant of MolJO with beam search width 10, selecting a tenth of top scoring molecules and showing that it is more effective than oracle-based methods once in a similar setting. Moreover, the higher diversity of DecompOpt and MolJO suggests the superiority of 3D structure-aware generative models over 2D optimization of RGA (row 14 vs. row 10 & 11).
For gradient-based method TAGMol (Dorna et al., 2024), although it produces seemingly promising high affinity binders, they come at the expense of sacrificed molecular properties like QED and SA, demonstrating the suboptimal control of coordinate-only guidance signals. Notably, the ratio of all-better samples is below 17% for all other baselines, and MolJO is twice as effective (39.8%) in generating feasible drug candidates that pass this criteria, as shown in Figure 1.
Table 2: Properties of molecules with a larger average size, where Vina stands for Vina Dock Avg., SR for Success Rate. Top-1 results are highlighted with bold text.
Methods Vina QED SA SR Size
Reference -7.45 0.48 0.73 25.0% 22.8
DecompDiff -8.39 0.45 0.61 24.5% 29.4
DecompOpt -9.01 0.48 0.65 52.5% 32.9
MolCRAFT -9.25 0.46 0.62 36.6% 29.4
MolJO -10.53 0.50 0.72 64.2% 30.0
MolJO excels even in optimizing large OOD molecules.
Note that for fair comparison, we restrict the size of generated molecules by reference molecules so that both generative models and optimization methods navigate the similar chemical space, as we observe a clear correlation between properties and sizes in Figure 6. For model variants capable of exploring larger number of atoms, we report the results in Table 2 with sizes, where MolJO consistently outperforms other baselines, demonstrating its robustness. A detailed discussion can be found in Appendix E.
5.3 Constrained Optimization
Constrained optimization seeks to optimize the input reference molecules for enhanced properties while retaining specific structures. We generalize our framework with such structural control and show its potential for pharmaceutical use cases including R-group optimization and scaffold hopping, achieved by infilling. Details of this task are in Appendix D.2.
Table 3: Constrained optimization results, where Redesign means R-group optimization with fragments of the same size redesigned, Growing means fragment growing into larger size, Hopping means scaffold hopping. (↑) / (↓) indicates a larger / smaller number is better. Top-1 highlighted in bold.
@@ 5.3 Constrained Optimization
The visualization for constrained optimization is shown in Figure 4. It can be seen that the optimized molecules establish more key interactions with the protein pockets, thus binding more tightly to the active sites. For example, the optimized molecule for 2PC8 retains the key interaction formed by its scaffold, with R-group grown deeper inside the pocket, forming another two $\pi$-$\pi$ stackings.
@@ 5.4 Ablation Studies
We conduct ablation studies to thoroughly validate our design. More details are left in Appendix F.2.
Figure 5: Ablation studies of joint optimization over atom types and coordinates, where w/o type means the gradient is disabled for types, w/o coord disables the gradient for coordinates.
@@
To validate our choice of joint guidance over different modalities, we ablate the gradient for coordinates or types. As shown in Figure 5, utilizing gradients to guide both data modalities is consistently better than applying single-modality gradient only. For affinities, optimizing coordinates is effective in improving the spatial interactions, while for drug-like properties, guidance over atom types plays a crucial role. This underscores the significance of deriving appropriate guidance form jointly, and again supports our finding that a single coordinate guidance as in TAGMol is insufficient and yields suboptimal results.
@@
We denote sampling $\bm{\theta}_{i}$ according to Eq. 11 Vanilla for sampling only one $y\sim p_{S}$ in Eq. 1, Vanilla MC for a Monte Carlo estimate that integrates over multiple $y$, and sampling proposed by MolCRAFT is denoted Full B.C. as it corresponds to setting $k=n=200$ in Eq. 13, and for B.C. we set $k=130$ as backward correction steps. Table 4 shows that our method of keeping a history and selectively correcting the past not only improves the original unguided sampling, but also yields better results with guidance. Note that for vanilla case, the gradient guidance does not work as much probably due to the suboptimal history, while correcting a sufficient number of past steps boosts the optimization.
Table 4: Performances of no backward correction (Vanilla & Vanilla MC), fully corrected strategy (Full B.C.) and backward correction strategy (B.C.). Positive numbers in green show the relative improvement under guidance, while non-positive numbers in black indicate no performance gain.
Sampling Strategy Vina Score ($\downarrow$) Vina Min ($\downarrow$) QED ($\uparrow$) SA ($\uparrow$)
Avg. Med. Avg. Med. Avg. Avg.
w/o Guide Vanilla (k=1) -5.23 -5.81 -6.30 -6.17 0.46 0.62
Vanilla MC (k=1) -6.25 -6.70 -7.01 -7.05 0.51 0.61
Full B.C. (k=200) -6.22 -6.94 -7.14 -7.13 0.49 0.68
B.C. (k=130) -6.50 -7.00 -7.03 -7.14 0.49 0.69
w/ Guide Vanilla (k=1) -5.47 (+4.6%) -5.89 (+1.4%) -6.29 (-0.2%) -6.31 (+2.3%) 0.46 (+0.0%) 0.62 (+0.0%)
Vanilla MC (k=1) -6.62 (+5.9%) -7.27 (+8.5%) -7.74 (+10.4%) -7.79 (+10.5%) 0.55 (+7.8%) 0.65 (+6.6%)
Full B.C. (k=200) -7.42 (+19.3%) -7.98 (+15.0%) -8.25 (+15.5%) -8.24 (+15.6%) 0.54 (+10.2%) 0.76 (+11.8%)
B.C. (k=130) -7.52 (+15.7%) -8.06 (+15.1%) -8.34 (+18.6%) -8.40 (+17.6%) 0.56 (+14.3%) 0.77 (+11.6%)
@@ 6 Conclusion
We present MolJO, the first joint gradient-based SE(3)-equivariant framework to solve the structure-based molecule optimization problems, which only requires training energy functions as a proxy to predict molecular property, instead of expensive oracle simulations. The general framework further equips gradient-based optimization method with backward correction strategy, offering a flexible trade-off between exploration and exploitation. Experiments show that MolJO is able to improve the binding affinity of molecules by establishing more key interactions and enhance drug-likeness and synthesizability, achieving state-of-the-art performance when benchmarked on CrossDocked2020 (Success Rate 51.3%, Vina Dock -9.05 and SA 0.78), together with 4×4$\times$4 × improvement compared to gradient-based counterpart and 2×2$\times$2 × “Me-Better” Ratio as much as other 3D baselines.
Ethics Statement
@@
Reproducibility Statement
Reproducibility of this worked is ensured by detailed descriptions of experimental setup in Section 5.1 and implementation in Appendix D.
@@ Appendix A Overview of Bayesian Flow Networks
The generative process for the receiver begins with a prior distribution (referred to as input distribution $p_{I}(\mathbf{m}|\bm{\theta})=\prod_{d=1}^{N}p_{I}(m^{(d)}|\theta^{(d)})$ that is also factorized for $N$-dimensional data) that defines its initial belief about the data. For continuous data, the prior can be chosen as a Gaussian (Song et al., 2024; Qu et al., 2024), or other distribution such as von Mises distribution (Anonymous, 2024), while for discrete data, the prior is modeled as a uniform categorical distribution. Then, the receiver uses its belief $\bm{\theta}$ with the help of neural network to model the inter-dependency among dimensions and compute the output distribution $p_{O}(\hat{\mathbb{m}}|\bm{\theta},\mathbf{p};t)$, which represents its estimate of the possible reconstructions of original data $\mathbf{m}$, and is used to construct the receiver distribution (Eq. 3).
The Bayesian update function in Eq. 1 defines how the prior belief $\bm{\theta}_{0}$ is updated to the conjugate posterior $\bm{\theta}_{t}$. Ideally, the update requires aggregating all possible noisy latents $\mathbf{y}$ from the sender distribution $p_{S}$, while during the actual generative process, there is only receiver distribution $p_{R}$, leaving an exposure bias, and different approximations determine different forms of mapping $f$ to the posterior (Eq. 2), showcasing the flexibility in the design space of BFN.
Through the iterative communication between sender and receiver, the receiver progressively updates its belief of the underlying parameters, and training is achieved by minimizing the divergence between the sender and receiver distributions (Eq. 4). This is analogous to the way Bayesian inference works in parameter estimation: as more noisy data $\mathbf{y}$ is observed, the receiver’s posterior belief about the data $\mathbf{m}$ becomes increasingly accurate, which implies the reconstruction would be made easier.
@@ Appendix B Sender Distribution for Discrete Data
The continuous parameter $\mathbf{z}$ for discrete types $\mathbf{v}$ is updated by observed noisy $\mathbf{y}^{v}$. Here we briefly introduce how to configure the sender so that $\mathbf{y}^{v}$ follows a Gaussian as well. For detailed derivation, we refer the readers to Graves et al. (2023).
@@ C.1 Proof of Guided Bayesian Update Distribution
If a random vector $\mathbf{X}$ has probability density $f(\mathbf{x})\propto\mathcal{N}(\mathbf{x}|\bm{\mu},\bm{\Sigma})e^{\mathbf{c}^{\mathrm{T}}\mathbf{x}}$, where $\mathbf{c}$ is a constant vector with the same dimension as $\mathbf{X}$, then $\mathbf{X}\sim\mathcal{N}(\bm{\mu}+\bm{\Sigma}\mathbf{c},\bm{\Sigma})$.
@@
$\log f(\mathbf{x})=$ $C-\frac{1}{2}(\mathbf{x}-\bm{\mu})^{\mathrm{T}}\bm{\Sigma}^{-1}(\mathbf{x}-\bm{\mu})+\mathbf{c}^{\mathrm{T}}\mathbf{x}$
$=$ $C^{\prime}-\frac{1}{2}(\mathbf{x}-\bm{\mu}-\bm{\Sigma}\mathbf{c})^{\mathrm{T}}\bm{\Sigma}^{-1}(\mathbf{x}-\bm{\mu}-\bm{\Sigma}\mathbf{c})$ (64)
@@
Assuming $\bm{\mu}_{i}\sim\mathcal{N}(\bm{\mu}_{\phi},\sigma)$ and $\mathbf{y}_{i}\sim\mathcal{N}(\mathbf{y}_{\phi},\sigma^{\prime})$ in the original generative process of BFN, we can approximately sample $\bm{\mu}_{i},\mathbf{y}_{i}$ from the guided transition kernel $\pi(\bm{\theta}_{i}|\bm{\theta}_{i-1})$ according to Eq. 5 and 6.
@@
Under the definition of $\pi$, with the parameters $\bm{\theta}_{i}=(\bm{\mu}_{i},\mathbf{z}_{i})$ redefined as $(\bm{\mu}_{i},\mathbf{y}_{i})$, we have
$\pi(\bm{\mu}_{i},\mathbf{y}_{i}|\bm{\theta}_{i-1})\propto p_{\phi}(\bm{\mu}_{i},\mathbf{y}_{i}|\bm{\theta}_{i-1})p_{E}(\bm{\mu}_{i},\mathbf{y}_{i})$ (65)
@@
Eq. 14 and 15 guarantee that $\bm{\mu}_{i}\sim\mathcal{N}(\bm{\mu}_{\phi},\sigma)$, $\mathbf{y}_{i}\sim\mathcal{N}(\mathbf{y}_{\phi},\sigma^{\prime})$. Plugging $p_{E}(\bm{\mu}_{i},\mathbf{y}_{i})\propto e^{-E(\bm{\mu}_{i},\mathbf{y}_{i},t_{i})}$ into Eq. 65, we get
$\pi(\bm{\mu}_{i},\mathbf{y}_{i}|\bm{\theta}_{i-1})\propto\mathcal{N}\left(\bm{\mu}_{\phi},\sigma\right)\mathcal{N}\left(\mathbf{y}_{i}|\mathbf{y}_{\phi},\sigma^{\prime}\right)e^{-E(\bm{\mu}_{i},\mathbf{y}_{i},t_{i})}$ (66)
With $t_{i}$ fixed, perform a first-order Taylor expansion to $E(\bm{\mu},\mathbf{y},t_{i})$ at $(\bm{\mu}_{i-1},\mathbf{y}_{i-1})$:
$E(\bm{\mu}_{i},\mathbf{y}_{i},t_{i})\approx E(\bm{\mu}_{i-1},\mathbf{y}_{i-1},t_{i})-\mathbf{g}_{\bm{\mu}}^{\mathrm{T}}(\bm{\mu}_{i}-\bm{\mu}_{i-1})-\mathbf{g}_{\mathbf{y}}^{\mathrm{T}}(\mathbf{y}_{i}-\mathbf{y}_{i-1})$ (67)
where gradient $\mathbf{g}_{\bm{\mu}}=-\nabla_{\bm{\mu}}E(\bm{\theta},t_{i})|_{\bm{\theta}=\bm{\theta}_{i-1}}$, $\mathbf{g}_{\mathbf{y}}=-\nabla_{\mathbf{y}}E(\bm{\theta},t_{i})|_{\bm{\theta}=\bm{\theta}_{i-1}}$. Substitute it into Eq. 66:
$\pi(\bm{\mu}_{i},\mathbf{y}_{i}|\bm{\theta}_{i-1})\stackrel{{\scriptstyle\text{apx}}}{{\propto}}\mathcal{N}\left(\bm{\mu}_{i}|\bm{\mu}_{\phi},\sigma\right)\mathcal{N}\left(\mathbf{y}_{i}|\mathbf{y}_{\phi},\sigma^{\prime}\right)$ $e^{\mathbf{g}_{\bm{\mu}}^{\mathrm{T}}\bm{\mu}_{i}+\mathbf{g}_{\mathbf{y}}^{\mathrm{T}}\mathbf{y}_{i}}$ (68)

### v3

@@ 5.2 Unconstrained Optimization
[…]
Table 2: Constrained optimization results, where Redesign means R-group optimization with fragments of the same size redesigned, Growing means fragment growing into larger size, Hopping means scaffold hopping.
@@
The optimized distribution greatly improves upon the original generated distribution, as shown in Table 1 (row 14 vs. row 9).
As shown in Table 1, our model achieves state-of-the-art in affinity-related metrics while being highly drug-like, with the best Success Rate of 51.3%, a four-fold improvement over TAGMol (row 14 vs. row 13).
RGA (Fu et al., 2022) and DecompOpt (Zhou et al., 2024a) show satisfactory Success Rate, enjoying the advantage of oracle-based screening at some expense of diversity, while AutoGrow4 (Spiegel & Durrant, 2020) falls short in QED, yielding a suboptimal Success Rate. Given the same concentration use of Z-score (Zhou et al., 2024a), we report a variant of MolJO with top-of-$N$, selecting a tenth of top scoring molecules and showing that it is more effective than oracle-based methods once in a similar setting. Moreover, the higher diversity of DecompOpt and MolJO suggests the superiority of 3D structure-aware generative models over 2D optimization baselines (row 15 vs. row 10-12).
For gradient-based method TAGMol (Dorna et al., 2024), although it produces seemingly promising high affinity binders, they come at the expense of sacrificed molecular properties like QED and SA, demonstrating the suboptimal control of coordinate-only guidance signals. Notably, the ratio of all-better samples is below 17% for all other baselines, and MolJO is twice as effective (39.8%) in generating feasible drug candidates that pass this criteria (Fig. 1).
MolJO excels even in optimizing large OOD molecules.
Note that for fair comparison, we restrict the size of generated molecules by reference molecules so that both generative models and optimization methods navigate the similar chemical space, as we observe a clear correlation between properties and sizes in Fig. 4. For model variants capable of exploring larger number of atoms, we report the results in Table 3 with sizes, where MolJO consistently outperforms other baselines, demonstrating its robustness. A detailed discussion can be found in Appendix E.
Table 3: Properties of molecules with a larger average size, where Vina stands for Vina Dock Avg., SR for Success Rate.
Methods Vina QED SA SR Size
Reference -7.45 0.48 0.73 25.0% 22.8
DecompDiff -8.39 0.45 0.61 24.5% 29.4
DecompOpt -9.01 0.48 0.65 52.5% 32.9
MolCRAFT -9.25 0.46 0.62 36.6% 29.4
MolJO -10.53 0.50 0.72 64.2% 30.0
5.3 Constrained Optimization
Constrained optimization seeks to optimize the input reference molecules for enhanced properties while retaining specific structures. We generalize our framework with such structural control and show its potential for pharmaceutical use cases including R-group optimization and scaffold hopping, achieved by infilling (details in Appendix D.2).
@@ 5.3 Constrained Optimization
Table 4: Performances of no correction (Vanilla), SDE and backward correction strategy (B.C.) without and with gradient guidance. Positive numbers in green show the relative improvement, while non-positive numbers in black indicate no performance gain.
Grad Sampling Vina Score ($\downarrow$) Vina Min ($\downarrow$) Vina Dock ($\downarrow$) QED ($\uparrow$) SA ($\uparrow$)
Avg. Med. Avg. Med. Avg. Med. Avg. Avg.
Vanilla -5.23 -5.81 -6.30 -6.17 -7.37 -7.31 0.46 0.62
SDE -6.62 -7.08 -7.31 -7.24 -8.22 -8.32 0.51 0.65
B.C. -6.50 -7.00 -7.03 -7.14 -7.95 -7.87 0.49 0.69
Vanilla -5.47 (+4.6%) -5.89 (+1.4%) -6.29 (-0.2%) -6.31 (+2.3%) -7.49 (+2.2%) -7.46 (+2.1%) 0.46 (+0.0%) 0.62 (+0.0%)
SDE -7.11 (+7.4%) -7.53 (+6.3%) -7.76 (+6.1%) -7.73 (+6.8%) -8.39 (+2.1%) -8.66 (+4.1%) 0.50 (-1.9%) 0.68 (+4.6%)
B.C. -7.52 (+15.7%) -8.06 (+15.1%) -8.34 (+18.6%) -8.40 (+17.6%) -9.11 (+14.6%) -9.25 (+17.5%) 0.56 (+14.3%) 0.77 (+11.6%)
@@
The visualization for constrained optimization is shown in Fig. 3. It can be seen that the optimized molecules establish more key interactions with the protein pockets, thus binding more tightly to the active sites. For example, the optimized molecule for 2PC8 retains the key interaction formed by its scaffold, with R-group grown deeper inside the pocket, forming another two $\pi$-$\pi$ stackings.
@@ 5.4 Ablation Studies
We conduct ablation studies to thoroughly validate our design. More details are left to Appendix F.2. For all the 100 test proteins, we sample 10 molecules each.
@@
To validate our choice of joint guidance over different modalities, we ablate the gradient for coordinates or types. As shown in Table 11, utilizing gradients to guide both data modalities is consistently better than applying single-modality gradient only. For affinities, optimizing coordinates is effective in improving the spatial interactions, while for drug-like properties, guidance over atom types plays a crucial role. This underscores the significance of deriving appropriate guidance form jointly, and again supports our finding that a single coordinate guidance as in TAGMol is insufficient and yields suboptimal results.
@@
We denote sampling $\bm{\theta}_{i}$ according to Eq. 10 Vanilla for point estimate of $y\sim p_{S}$ in Eq. 1, advanced SDE proposed by Xue et al. (2024) with classifier guidance, and for B.C. we set $k=130$ as backward correction steps. Table 4 shows that our method of correcting the past yields better results with guidance. Note that for vanilla case, the gradient guidance does not work as much probably due to the suboptimal history, and SDE-based classifier guidance may have suffered from discretization errors, while correcting a sufficient number of past steps shows consistent boosts.
@@ 6 Conclusion
We present MolJO, the joint gradient-based SE(3)-equivariant framework to solve the structure-based molecule optimization problems, which only requires differentiable energy functions instead of expensive oracle simulations. The general framework further equips gradient-based optimization method with backward correction strategy, offering a flexible trade-off between exploration and exploitation. Experiments show that MolJO is able to improve the binding affinity of molecules by establishing more key interactions and enhance drug-likeness and synthesizability, achieving state-of-the-art performance on CrossDocked2020 (Success Rate 51.3%, Vina Dock -9.05 and SA 0.78), together with 4×4$\times$4 × improvement compared to gradient-based counterpart and 2×2$\times$2 × “Me-Better” Ratio as other 3D baselines.
Impact Statement
@@
Acknowledgments
This work is supported the Natural Science Foundation of China (Grant No. 62376133) and sponsored by Beijing Nova Program (20240484682) and the Wuxi Research Institute of Applied Technologies, Tsinghua University (20242001120).
@@ Appendix A Overview of Bayesian Flow Networks
The generative process for the receiver begins with a prior distribution (referred to as input distribution $p_{I}(\mathbf{m}|\bm{\theta})=\prod_{d=1}^{N}p_{I}(m^{(d)}|\theta^{(d)})$ that is also factorized for $N$-dimensional data) that defines its initial belief about the data. For continuous data, the prior can be chosen as a Gaussian (Song et al., 2024; Qu et al., 2024), or other distribution such as von Mises distribution (Wu et al., 2025), while for discrete data, the prior is modeled as a uniform categorical distribution. Then, the receiver uses its belief $\bm{\theta}$ with the help of neural network to model the inter-dependency among dimensions and compute the output distribution $p_{O}(\hat{\mathbb{m}}|\bm{\theta},\mathbf{p};t)$, which represents its estimate of the possible reconstructions of original data $\mathbf{m}$, and is used to construct the receiver distribution (Eq. 2).
The Bayesian update function in Eq. 1 defines how the prior belief $\bm{\theta}_{0}$ is updated to the conjugate posterior $\bm{\theta}_{t}$. Ideally, the update requires aggregating all possible noisy latents $\mathbf{y}$ from the sender distribution $p_{S}$, while during the actual generative process, there is only receiver distribution $p_{R}$, leaving an exposure bias, and different approximations determine different forms of mapping $\theta=f(\mathbf{y})$ to the posterior, showcasing the flexibility in the design space of BFN.
Through the iterative communication between sender and receiver, the receiver progressively updates its belief of the underlying parameters, and training is achieved by minimizing the divergence between the sender and receiver distributions (Eq. 3). This is analogous to the way Bayesian inference works in parameter estimation: as more noisy data $\mathbf{y}$ is observed, the receiver’s posterior belief about the data $\mathbf{m}$ becomes increasingly accurate, which implies the reconstruction would be made easier. It is shown that BFN for Gaussian priors can be sampled from the view of SDE (Xue et al., 2024), but this kind of sampling displays distinct behaviors empirically, possibly due to discretization errors.
@@ Appendix B Sender Distribution for Discrete Data
The continuous parameter $\bm{\theta}^{v}$ for discrete types $\mathbf{v}$ is updated by observed noisy $\mathbf{y}^{v}$. Here we briefly introduce how to configure the sender so that $\mathbf{y}^{v}$ follows a Gaussian as well. For detailed derivation, we refer the readers to Graves et al. (2023).
@@ C.1 Proof of Guided Bayesian Update Distribution
If a random vector $\mathbf{X}$ has probability density $f(\mathbf{x})\propto\mathcal{N}(\mathbf{x}|\bm{\theta}^{x},\bm{\Sigma})e^{\mathbf{c}^{\mathrm{T}}\mathbf{x}}$, where $\mathbf{c}$ is a constant vector with the same dimension as $\mathbf{X}$, then $\mathbf{X}\sim\mathcal{N}(\bm{\theta}^{x}+\bm{\Sigma}\mathbf{c},\bm{\Sigma})$.
@@
$\log f(\mathbf{x})=$ $C-\frac{1}{2}(\mathbf{x}-\bm{\theta}^{x})^{\mathrm{T}}\bm{\Sigma}^{-1}(\mathbf{x}-\bm{\theta}^{x})+\mathbf{c}^{\mathrm{T}}\mathbf{x}$
$=$ $C^{\prime}-\frac{1}{2}(\mathbf{x}-\bm{\theta}^{x}-\bm{\Sigma}\mathbf{c})^{\mathrm{T}}\bm{\Sigma}^{-1}(\mathbf{x}-\bm{\theta}^{x}-\bm{\Sigma}\mathbf{c})$ (15)
@@
Assuming $\bm{\theta}^{x}_{i}\sim\mathcal{N}(\bm{\theta}^{x}_{\phi},\sigma)$ and $\mathbf{y}^{v}_{i}\sim\mathcal{N}(\mathbf{y}_{\phi},\sigma)$ by definition of the generative process of BFN, we can approximately sample $\bm{\theta}^{x}_{i},\mathbf{y}^{v}_{i}$ from the guided transition kernel $\pi(\bm{\theta}_{i}|\bm{\theta}_{i-1})$ according to Eq. 4 and 5.
@@
Under the definition of $\pi$, with the parameters $\bm{\theta}_{i}=[\bm{\theta}^{x}_{i},\bm{\theta}^{v}_{i}]$ referred to as $[\bm{\theta}^{x}_{i},\bm{\theta}^{v}_{i}(\mathbf{y}^{v}_{i})]$ and a slight abuse of notation, we have
$\pi(\bm{\theta}^{x}_{i},\mathbf{y}^{v}_{i}|\bm{\theta}_{i-1})\propto p_{\phi}(\bm{\theta}^{x}_{i},\mathbf{y}^{v}_{i}|\bm{\theta}_{i-1})p_{E}(\bm{\theta}^{x}_{i},\mathbf{y}^{v}_{i})$ (16)
@@
Eq. 13 and 4.2 guarantee that $\bm{\theta}^{x}_{i}\sim\mathcal{N}(\bm{\theta}^{x}_{\phi},\sigma^{x})$, $\mathbf{y}^{v}_{i}\sim\mathcal{N}(\mathbf{y}_{\phi},\sigma^{v})$. Plugging $p_{E}(\bm{\theta}^{x}_{i},\mathbf{y}^{v}_{i})\propto e^{-E(\bm{\theta}^{x}_{i},\mathbf{y}^{v}_{i},t_{i})}$ into Eq. 16, we get
$\pi(\bm{\theta}^{x}_{i},\mathbf{y}^{v}_{i}|\bm{\theta}_{i-1})\propto\mathcal{N}\left(\bm{\theta}^{x}_{\phi},\sigma^{x}\right)\mathcal{N}\left(\mathbf{y}^{v}_{i}|\mathbf{y}_{\phi},\sigma^{v}\right)e^{-E(\bm{\theta}^{x}_{i},\mathbf{y}^{v}_{i},t_{i})}$ (17)
With $t_{i}$ fixed, perform a first-order Taylor expansion to $E(\bm{\theta}^{x},\mathbf{y}^{v},t_{i})$ at $(\bm{\theta}^{x}_{i-1},\mathbf{y}^{v}_{i-1})$:
$E(\bm{\theta}^{x}_{i},\mathbf{y}^{v}_{i},t_{i})\approx E(\bm{\theta}^{x}_{i-1},\mathbf{y}^{v}_{i-1},t_{i})-\mathbf{g}_{\bm{\theta}^{x}}^{\mathrm{T}}(\bm{\theta}^{x}_{i}-\bm{\theta}^{x}_{i-1})-\mathbf{g}_{\mathbf{y}}^{\mathrm{T}}(\mathbf{y}^{v}_{i}-\mathbf{y}^{v}_{i-1})$ (18)
where gradient $\mathbf{g}_{\bm{\theta}^{x}}=-\nabla_{\bm{\theta}^{x}}E(\bm{\theta},t_{i})|_{\bm{\theta}=\bm{\theta}_{i-1}}$, $\mathbf{g}_{\mathbf{y}}=-\nabla_{\mathbf{y}}E(\bm{\theta},t_{i})|_{\bm{\theta}=\bm{\theta}_{i-1}}$. Substitute it into Eq. 17:
$\pi(\bm{\theta}^{x}_{i},\mathbf{y}^{v}_{i}|\bm{\theta}_{i-1})\stackrel{{\scriptstyle\text{apx}}}{{\propto}}\mathcal{N}\left(\bm{\theta}^{x}_{i}|\bm{\theta}^{x}_{\phi},\sigma^{x}\right)\mathcal{N}\left(\mathbf{y}^{v}_{i}|\mathbf{y}_{\phi},\sigma^{v}\right)$ $e^{\mathbf{g}_{\bm{\theta}^{x}}^{\mathrm{T}}\bm{\theta}^{x}_{i}+\mathbf{g}_{\mathbf{y}}^{\mathrm{T}}\mathbf{y}^{v}_{i}}$ (19)